Given non-negative integers such that (108^a) ⋅ (288^b) ⋅ (36^c) divided by 6^220. Find the smallest possible value of a + b + c.
1
Expert's answer
2020-12-15T13:40:36-0500

"108=2^2\cdot 3^3"

"288=2^5\cdot 3^2"

"36=2^2\cdot 3^2"

So "108^a\cdot 288^b\cdot 36^c=2^{2a+5b+2c}\cdot 3^{3a+2b+2c}"

"6^{220}=2^{220}\cdot 3^{220}"

So "2a+5b+2c\ge 220" and "3a+2b+2c\ge 220"

Denote "a+b+c" by "u", then we have "2u+3b\ge 220" and "2u+a\ge 220". Since "c\ge 0", we also have "u\ge a+b".

"u=80,a=60,b=20" (then "c=0") satisfied these 3 inequalities.

If we take "u<80", then "b\ge\frac{220-2u}{3}>20" and "a\ge 220-2u>60", and we obtain "a+b>60+20=80>u". It is impossible, so smallest value of "u" is "80", when "a=60, b=20, c=0"

Answer: "80"